The average mass of 120 goats is 12.5kg.if two goats are added, the average mass is 12.6kg.what is the mass of the extra goats.
1
Expert's answer
2020-02-24T11:30:39-0500

"\frac{120*12.5+x}{120+2}=12.6"

"x=12.6(120+2)-120*12.5=37.2"

The mass of two extra goats is 37.2 kg.